OpenStudy (anonymous):

Movies often use models to create special effects. In the movie Son of Godzilla, Godzilla appears to be 67 feet tall and was played by a man in a 7.5 ft rubber suit. Godzilla’s son appeared to be 45 feet tall in the movie; how tall was the rubber suit for the actor playing the son? (Round to the nearest hundredth)(Points : 2) 4.86 ft 5.04 ft 5.87 ft 6.14 ft

OpenStudy (anonymous):

7.5/67 = x/45 67x=337.5 x=5.04